OpenStudy (anonymous):

What are the solutions of the system? y=x^2-3x+2 y=4x-4 A. (1, 0) and (–6, –28) B. (1, 20) and (6, 0) C. (1, 0) and (6, 20) D. no solution

OpenStudy (phi):

i would try each case. (1,0) means x=1 and y=0. Does this pair work for both equations?

OpenStudy (anonymous):

Try setting the two equal to each other first, that will tell you when they intersect. :-) \[x^2-3x+2=4x-4\] \[x^2-7x+6=0\] \[(x-6)(x-1)=0\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then plug in these two x solutions as inputs into either one of the equations, doesn't matter which as they'll both give the same thing. That give you your y coordinate output at that x coordinate input, the very definition of what a function is btw.

OpenStudy (anonymous):

that means the answer is C

OpenStudy (anonymous):

\[y=f(x), f(0)=4x-4=4(1)-4=0\]

OpenStudy (anonymous):

like that, if it wasn't clear :-)

OpenStudy (anonymous):

so is it C

OpenStudy (anonymous):

\[f(6)=4x-4=4(6)-4=24-4=20\]

OpenStudy (anonymous):

Try the other equation if you have doubt, it gives you 20 out when x is 6 as well.

OpenStudy (anonymous):

thanks might need more help soon